OpenStudy (anonymous):

Prove that the greatest area of a rectangle inscribed in a triangle is one half of the area of the triangle. Use calculus

OpenStudy (amistre64):

a triangle can be defined as a line: y = mx + b

OpenStudy (amistre64):

the area of a rectangle from this setup is 2xy

OpenStudy (anonymous):

ok I get the triangle being defined as a line part but wouldn't the area of a rectangle simply be xy?

OpenStudy (amistre64):

using lagrange multipliers .... f(x,y) = 2xy ; g(x,y) = mx - y + b fx = 2y ; Lgx = mL ; y = mL/2 fy = x ; Lgx = -L ; x = -L plugging these into the function g: y-mx = b mL/2 - m(-L) = b solve for L 3mL/2 = b L = 2b/3m; x= -2b/3m, y = -1b/3

OpenStudy (amistre64):

since the line only represents half the triangle .... height as y and width as 2x

OpenStudy (amistre64):

if we want to make it general, we would need to be able to define all triangles .... at the moment this is just for right/isosolese :/

OpenStudy (anonymous):

well I think the problem is I haven't learned lagrange variables

OpenStudy (anonymous):

sorry lagrange multipliers

OpenStudy (amistre64):

lagrange is a method ... im sure there are other methods just as suitable, i just cant think what theyd be at the moment :)

OpenStudy (anonymous):

okay... im sorry but I really didn't get your process

OpenStudy (amistre64):